OpenStudy (anonymous):

natasha rides her bike (at a constant speed) for 5 hours with help by a wind of 2 mph. pedaling at the same rate, the trip back against the wind takes 9 hours. Find the total roundtrip distance she traveled.

OpenStudy (kinggeorge):

For this, you need to remember the equation distance=rate x time. When she was riding with the help of the wind, her rate was \(r+2\) mph, where \(r\) would be her rate in no wind. Against the wind, her rate would be \(r-2\). We also know that the distance both with and against the wind was the same. So we have two equations.\[d=(r+2)*5\]\[d=(r-2)*9.\]Does this make sense so far?

OpenStudy (anonymous):

yes i have that down alread

OpenStudy (anonymous):

y

OpenStudy (anonymous):

i just dont know what to do next

OpenStudy (kinggeorge):

Next we set the two equations equal to each other.\[(r+2)*5=(r-2)*9\implies 5r+10=9r-18\implies28=4r\implies r=7.\]So the rate Natasha would go if there were no wind, is 7mph. Now we can just plug this value back into one of the original equations for \(d\).\[d=(r+2)*5=(7+2)*5=9*5=45.\]Did this clear things up for you?

OpenStudy (kinggeorge):

However, you need to remember that this \(d\) is just one way. So the roundtrip distance would be \(2d=90\).

OpenStudy (anonymous):

Wow thank you so much. I appreciate it!

OpenStudy (kinggeorge):

You're very welcome.